# Many-body Fu-Kane-Mele index

### v1

@@ B.1 Locality estimates for Fermi projections
We take $c,\xi>0$ such that $\left|{\mathscr{K}}_{\phi}^{h}(x,i;y,j)\right|\leq c{\mathrm{e}}^{-2\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (cf. Lemma B.1) and
$\kappa=2\sup_{x_{1}\in{\mathbb{Z}}}\sum_{z\in{\mathbb{Z}}^{2}}\,{\mathrm{e}}^{-\xi\left|z_{1}-x_{1}\right|-3\xi\left|z_{2}\right|}.$ (B.9)
Then the claimed bound on $a_{n}(x,y)$ holds for $n=1$ by that same Lemma. We proceed by induction using the again second claim of Lemma B.1 to obtain
$a_{n}(x,y)$ $\leq\sup_{\phi\in[-\pi,\pi]}\sup_{i\in\{1,2\}}\sum_{z,i^{\prime}}\left|{\mathscr{K}}_{\phi}(x,i;z,i^{\prime})\right|\times a_{n-1}(z,y)$ (B.10)
$\leq c^{n}\kappa^{n-2}\sum_{z,i^{\prime}}\,{\mathrm{e}}^{-2\xi(\left|x_{1}-z_{1}\right|+\left|x_{2}\right|+\left|z_{2}\right|)}\times{\mathrm{e}}^{-\xi(\left|z_{1}-y_{1}\right|+\left|z_{2}\right|+\left|y_{2}\right|)}$ (B.11)
noting that ${\mathrm{e}}^{-\xi\left|z_{1}-y_{1}\right|}\,{\mathrm{e}}^{-2\xi\left|x_{1}-z_{1}\right|}\leq{\mathrm{e}}^{-\xi\left|x_{1}-y_{1}\right|}\,{\mathrm{e}}^{-\xi\left|x_{1}-z_{1}\right|}$ this is easily bounded by
$\leq c^{n}\kappa^{n-1}{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (B.12)
which proves the bound on $a_{n}$ for all $n$. The bound on $b_{n}$ is obtained in exactly the same way and recalling that ${\mathscr{K}}_{\phi}=\Pi_{L}{\mathscr{K}}_{\phi}^{h}\Pi_{L}$. ∎
Recall that ${\mathscr{U}}_{\phi}$ is the unique solution to ${\mathrm{i}}\big{(}{\mathrm{d}}{\mathscr{U}}_{\phi}/{\mathrm{d}}\phi\big{)}={\mathscr{K}}_{\phi}{\mathscr{U}}_{\phi}$ with initial condition ${\mathscr{U}}_{0}=\operatorname{\mathbbm{1}}$. Similarly we let ${\mathscr{U}}_{\phi}^{h}$ be the unique solution to ${\mathrm{i}}({\mathrm{d}}{\mathscr{U}}_{\phi}^{h}/{\mathrm{d}}\phi)={\mathscr{K}}^{h}_{\phi}{\mathscr{U}}_{\phi}^{h}$ with initial condition ${\mathscr{U}}^{h}_{0}=\operatorname{\mathbbm{1}}$.
Lemma B.3.
@@
$\left|({\mathscr{U}}_{\phi}^{h}-\operatorname{\mathbbm{1}})(x,i;y,j)\right|$ $\leq c{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (B.13)
$\left|({\mathscr{U}}_{\phi}-\operatorname{\mathbbm{1}})(x,i;y,j)\right|$ $\leq c{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}\chi_{\leq 0}(x_{1})\chi_{\leq 0}(x_{2})$ (B.14)
for all $x=(x_{1},x_{2}),y=(y_{1},y_{2})\in{\mathbb{Z}}^{2}$, all $i,j\in\{1,2\}$, and all $\phi\in[-\pi,\pi]$.
@@
We have
$\left|({\mathscr{U}}_{\phi}^{h}-\operatorname{\mathbbm{1}})(x,i;y,j)\right|\leq\sum_{n=1}^{\infty}\frac{\left|\phi\right|^{n}}{n!}a_{n}(x,y)$ (B.15)
where $a_{n}(x,y)$ was defined in (B.5). The first claim of the Lemma now follows straightforwardly from the bound on $a_{n}(x,y)$ provided by Lemma B.2. The second claim follows in the same way using $b_{n}(x,y)$ instead. ∎
${\mathscr{U}}_{\phi}$ is a restriction of ${\mathscr{U}}_{\phi}^{h}$ to the left horizontal axis. Since ${\mathscr{U}}_{\phi}^{h}$ acts non-trivially only near the horizontal axis so we expect ${\mathscr{U}}^{h}_{\phi}-{\mathscr{U}}_{\phi}$ to be localised near the right horizontal axis.
Let us introduce a new family of projections
${\mathscr{Q}}_{\phi}:={\mathrm{e}}^{{\mathrm{i}}(\phi+\pi)\Pi_{h}}{\mathscr{P}}_{-}{\mathrm{e}}^{-{\mathrm{i}}(\phi+\pi)\Pi_{h}}$ (B.16)
@@
which satisfies $\partial_{\phi}{\mathscr{Q}}_{\phi}=-{\mathrm{i}}[{\mathscr{L}}_{\phi},{\mathscr{Q}}_{\phi}]$ with ${\mathscr{L}}_{\phi}={\mathrm{i}}[\partial_{\phi}{\mathscr{Q}}_{\phi},{\mathscr{Q}}_{\phi}]$. We let ${\mathscr{V}}_{(\phi_{2},\phi_{1})}$ be the unique unitaries such that ${\mathrm{i}}({\mathrm{d}}{\mathscr{V}}_{(\phi_{2},\phi_{1})}/{\mathrm{d}}\phi_{2})={\mathscr{L}}_{\phi_{2}}{\mathscr{V}}_{(\phi_{2},\phi_{1})}$ and ${\mathscr{V}}_{(\phi_{1},\phi_{1})}=\operatorname{\mathbbm{1}}$, then ${\mathscr{Q}}_{\phi_{2}}={\mathscr{V}}_{(\phi_{2},\phi_{1})}{\mathscr{Q}}_{\phi_{1}}{\mathscr{V}}_{(\phi_{1},\phi_{2})}$.
We want to compare the unitaries ${\mathscr{V}}_{(\phi_{1},\phi_{2})}$ generated by the ${\mathscr{L}}_{\phi}$ with the unitaries ${\mathscr{U}}_{\phi}^{h}$ generated by the ${\mathscr{K}}^{h}_{\phi}$. It is therefore convenient to extend the notation by letting ${\mathscr{U}}^{h}_{(\phi_{2},\phi_{1})}$ be the solution to ${\mathrm{i}}({\mathrm{d}}{\mathscr{U}}^{h}_{(\phi_{2},\phi_{1})}/{\mathrm{d}}\phi_{2})={\mathscr{K}}^{h}_{\phi_{2}}{\mathscr{U}}^{h}_{(\phi_{2},\phi_{1})}$ and ${\mathscr{U}}_{(\phi_{1},\phi_{1})}=\operatorname{\mathbbm{1}}$. Then ${\mathscr{U}}^{h}_{\phi}={\mathscr{U}}^{h}_{(\phi,0)}$.
By construction, the projections ${\mathscr{Q}}_{\phi}$ differ from ${\mathscr{P}}_{\phi}^{h}$ only near the right horizontal axis, and so the same holds for the generators ${\mathscr{L}}_{\phi}$ and ${\mathscr{K}}^{h}_{\phi}$ and for the unitaries ${\mathscr{V}}_{\phi_{2},\phi_{1}}$ and ${\mathscr{U}}^{h}(\phi_{2},\phi_{1})$.
Lemma B.4.
There are constants $c,\xi>0$ such that
$\left|({\mathscr{U}}^{h}_{\phi}-{\mathscr{U}}_{\phi})(x,i;y,j)\right|$ $\leq c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (B.17)
$\left|({\mathscr{Q}}_{\phi}-{\mathscr{P}}^{h}_{\phi})(x,i;y,j)\right|$ $\leq c\,{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (B.18)
$\left|({\mathscr{L}}_{\phi}-{\mathscr{K}}^{h}_{\phi})(x,i;y,j)\right|$ $\leq c\,{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (B.19)
$\left|({\mathscr{V}}_{(\phi_{2},\phi_{1})}-{\mathscr{U}}^{h}_{(\phi_{2},\phi_{1})})(x,i;y,j)\right|$ $\leq c\,{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (B.20)
for all $x=(x_{1},x_{2}),y=(y_{1},y_{2})\in{\mathbb{Z}}^{2}$, all $i,j\in\{1,2\}$, and all $\phi,\phi_{1},\phi_{2}\in[-\pi,\pi]$.
@@
The first bound : If $x,y\in R$ this follows immediately from Lemma B.3. To deal with the other cases, note first that
$\left|({\mathscr{U}}^{h}_{\phi}-{\mathscr{U}}_{\phi})(x,i;y,j)\right|\leq\sum_{n\geq 1}\,\frac{\left|\phi\right|^{n}}{n!}c_{n}(x,y)$ (B.21)
with
$c_{n}(x,y)=\sup_{\begin{subarray}{c}\vec{\phi}\in[-\pi,\pi]^{n}\\ i,j\in\{1,2\}\end{subarray}}\,\left|({\mathscr{K}}^{h}_{\phi_{1}}\cdots{\mathscr{K}}^{h}_{\phi_{n}})(x,i;y,j)-({\mathscr{K}}_{\phi_{1}}\cdots{\mathscr{K}}_{\phi_{n}})(x,i;y,j)\right|.$ (B.22)
For arbitrary $x\in L$ and $y\in{\mathbb{Z}}^{2}$ we find
$c_{n}(x,y)$ $=\sup_{\begin{subarray}{c}\vec{\phi}\in[-\pi,\pi]^{n}\\ i,j\in\{1,2\}\end{subarray}}\left|\sum_{\begin{subarray}{c}\vec{A}\in\{L,R\}^{n-1}\\ \text{not all }L\end{subarray}}({\mathscr{K}}^{h}_{\phi_{1}}\Pi_{A_{1}}\cdots\Pi_{A_{n-1}}{\mathscr{K}}_{\phi_{n}}^{h})(x,i;y,j)\right|$ (B.23)
noting that there are $12^{n}-$ 1 terms in the sum, that each summand contains at least one projection $\Pi_{R}$, and using Lemma B.2,
$\leq 2^{n}\sup_{k+l=n}\,\sum_{z\in R}\,a_{k}(x,z)a_{l}(z,y)\leq(2c)^{n}\kappa^{n-2}\sum_{z\in R}\,{\mathrm{e}}^{-\xi(\left|x_{1}-z_{1}\right|+\left|z_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|+2\left|z_{2}\right|)}$ (B.24)
since $\left|x_{1}-z_{1}\right|=\left|x_{1}\right|+\left|z_{1}\right|$, using the triangle inequality and performing the sum over $z$ there is a constant $D>0$ such that
$\leq D^{n}{\mathrm{e}}^{-\frac{\xi}{2}(\|x-y\|_{1}+\left|x_{1}\right|+\left|x_{2}\right|)}.$ (B.25)
It follows that
$\left|({\mathscr{U}}^{h}_{\phi}-{\mathscr{U}}_{\phi})(x,i;y,j)\right|\leq({\mathrm{e}}^{2\pi D}-1)\,{\mathrm{e}}^{-\frac{\xi}{2}(\|x-y\|_{1}+\left|x_{1}\right|+\left|x_{2}\right|)}$ (B.26)
which proves the claim for $x\in L$ and $y\in{\mathbb{Z}}$ arbitrary. The case where $x$ is arbitrary and $y\in L$ follows in the same way.
The second bound : We have
${\mathscr{Q}}_{\phi}-{\mathscr{P}}^{h}_{\phi}={\mathrm{e}}^{{\mathrm{i}}(\phi+\pi)\Pi_{h}}{\mathscr{U}}_{-\pi}\,{\mathscr{P}}\,{\mathscr{U}}_{-\pi}^{*}{\mathrm{e}}^{-{\mathrm{i}}(\phi+\pi)\Pi_{h}}-{\mathrm{e}}^{{\mathrm{i}}\phi\Pi_{h}}{\mathscr{P}}{\mathrm{e}}^{-{\mathrm{i}}\phi\Pi_{h}}$ (B.27)
so it is sufficient to establish the bound for ${\mathscr{U}}_{-\pi}\,{\mathscr{P}}\,{\mathscr{U}}_{-\pi}^{*}-{\mathrm{e}}^{-{\mathrm{i}}\pi\Pi_{h}}{\mathscr{P}}{\mathrm{e}}^{{\mathrm{i}}\pi\Pi_{h}}={\mathscr{U}}_{-\pi}\,{\mathscr{P}}\,{\mathscr{U}}_{-\pi}^{*}-{\mathscr{U}}^{h}_{-\pi}\,{\mathscr{P}}\,({\mathscr{U}}^{h}_{-\pi})^{*}$ which follows easily from the first bound.
The third bound : follows from
${\mathscr{L}}_{\phi}-{\mathscr{K}}_{\phi}^{h}=-[[\Pi_{h},{\mathscr{Q}}_{\phi}],({\mathscr{Q}}_{\phi}-{\mathscr{P}}_{\phi}^{h})]-[[\Pi_{h},({\mathscr{P}}_{\phi}^{h}-{\mathscr{Q}}_{\phi})],{\mathscr{P}}_{\phi}^{h}]$ (B.28)
and the second bound.
The fourth bound : follows straightforwardly from the analogous third bound on the generators. ∎
Proposition B.5.
The difference of projections ${\mathscr{P}}_{+}-{\mathscr{P}}_{-}$ is trace class.
@@
Fix $x,y\in{\mathbb{Z}}^{2}$ and $i,j\in\{1,2\}$. Since ${\mathscr{P}}_{+}-{\mathscr{P}}_{-}={\mathscr{U}}_{\pi}{\mathscr{P}}{\mathscr{U}}_{\pi}^{*}-{\mathscr{U}}_{-\pi}{\mathscr{P}}{\mathscr{U}}_{-\pi}^{*}$ we get from Lemma B.3 that
$\left|({\mathscr{P}}_{+}-{\mathscr{P}}_{-})(x,i;y,j)\right|\leq c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{L}))}$ (B.29)
for some constants $c,\xi>0$. This shows that ${\mathscr{P}}_{+}-{\mathscr{P}}_{-}$ is localized near the left horizontal axis, but it is also localized near the right horizontal axis. Indeed since ${\mathscr{P}}_{-}={\mathscr{Q}}_{\pi}={\mathscr{V}}_{(\pi,-\pi)}{\mathscr{Q}}_{-\pi}{\mathscr{V}}_{(\pi,-\pi)}^{*}={\mathscr{V}}_{(\pi,-\pi)}{\mathscr{P}}_{-}{\mathscr{V}}_{(\pi,-\pi)}^{*}$ we have ${\mathscr{P}}_{+}-{\mathscr{P}}_{-}={\mathscr{U}}_{(\pi,-\pi)}{\mathscr{P}}_{-}{\mathscr{U}}_{(\pi,-\pi)}^{*}-{\mathscr{V}}_{(\pi,-\pi)}{\mathscr{P}}_{-}{\mathscr{V}}_{(\pi,-\pi)}^{*}$ so using the fourth bound of Lemma B.4 we obtain
$\left|({\mathscr{P}}_{+}-{\mathscr{P}}_{-})(x,i;y,j)\right|\leq c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (B.30)
for some constants $c,\xi>0$.
Combining the bounds (B.29) and (B.30) we get
$\left|({\mathscr{P}}_{+}-{\mathscr{P}}_{-})(x,i;y,j)\right|\leq c{\mathrm{e}}^{-\frac{\xi}{2}(2\|x-y\|_{1}+\mathrm{dist}(x,l_{L})+\mathrm{dist}(x,l_{R}))}\leq c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\|x\|_{1}/2)}.$ (B.31)

### v2

@@ C.2 Comparison of quasi-adiabatic and spectral generators
By the gap assumption for ${\mathscr{H}}_{\phi}$ we can find a contour $\Gamma$ in the complex plane the circles the spectra of ${\mathscr{H}}_{\phi}$ and ${\mathscr{H}}_{\phi}^{h}$ below $\mu$ counterclockwise, and such that $\mathrm{dist}(\Gamma,\sigma({\mathscr{H}}_{\phi})),\mathrm{dist}(\Gamma,\sigma({\mathscr{H}}^{h}_{\phi}))$ are bounded below uniformly in $\phi\in[\phi^{(i)},\phi^{(f)}]$. Then we have
${\mathscr{P}}^{(\mu)}_{\phi}=-\frac{1}{2\pi{\mathrm{i}}}\oint_{\Gamma}{\mathrm{d}}z\,({\mathscr{H}}_{\phi}-z)^{-1},$ (C.12)
from which we obtain
$\partial_{\phi}{\mathscr{P}}_{\phi}^{(\mu)}=\frac{1}{2\pi{\mathrm{i}}}\oint_{\Gamma}{\mathrm{d}}z({\mathscr{H}}_{\phi}-z)^{-1}(\partial_{\phi}{\mathscr{H}}_{\phi})({\mathscr{H}}_{\phi}-z)^{-1}.$ (C.13)
The bounds (C.8) and (C.9) now follow from the exponential locality of the Hamiltonians ${\mathscr{H}}_{\phi},{\mathscr{H}}^{h}_{\phi}$, the Combes-Thomas estimates for the resolvents (see for example Theorem 10.5 of [2]) and the fact that $\partial_{\phi}{\mathscr{H}}_{\phi}$ is exponentially localized near the left half-line $l_{L}$ by construction. The bounds (C.10) follow in the same way using a Riesz projector formula for ${\mathscr{P}}_{\phi}^{h}$ and the fact that ${\mathscr{H}}^{h}_{\phi}-{\mathscr{H}}_{\phi}$ is exponentially localized near the right half-line $l_{R}$.
Finally, (C.11) follows from (C.8) and (C.9), and the definitions. ∎
Lemma C.3.
@@
$\left|({\mathscr{G}}^{(\mu)}_{\phi}-{\mathscr{K}}^{h}_{\phi})(x,i;y,j)\right|\leq c\,{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (C.14)
and
$\left|({\mathscr{G}}_{\phi}^{(\mu)}-{\mathscr{K}}_{\phi})(x,i;y,j)\right|\leq c\,{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\|x\|_{1})}$ (C.15)
for all $x,i,y,j$ and $\phi\in[\phi_{1},\phi_{2}]$.
@@
The first bound is an immediate consequence of the definitions ${\mathscr{G}}^{(\mu)}_{\phi}=-{\mathrm{i}}[\partial_{\phi}{\mathscr{P}}^{(\mu)}_{\phi},{\mathscr{P}}^{(\mu)}_{\phi}]$ and ${\mathscr{K}}^{h}_{\phi}=-{\mathrm{i}}[\partial_{\phi}{\mathscr{P}}^{h}_{\phi},{\mathscr{P}}^{h}_{\phi}]$ and the bounds provided by Lemma C.2.
Now recall that ${\mathscr{K}}_{\phi}=\Pi_{L}{\mathscr{K}}^{h}_{\phi}\Pi_{L}$ and combine this with Lemma C.1 to obtain $\left|{\mathscr{K}}_{\phi}(x,i;y,j)\right|\leq c^{\prime}{\mathrm{e}}^{-\xi^{\prime}(\|x-y\|_{1}+\mathrm{dist}(x,l_{L}))}$ for some $c^{\prime},\xi^{\prime}>0$. Together with the last bound of Lemma C.2 we obtain
$\left|({\mathscr{G}}_{\phi}^{(\mu)}-{\mathscr{K}}_{\phi})(x,i;y,j)\right|\leq c^{\prime\prime}{\mathrm{e}}^{-\xi^{\prime\prime}(\|x-y\|_{1}+\mathrm{dist}(x,l_{L}))}$ (C.16)
for some $c^{\prime\prime},\xi^{\prime\prime}>0$.
Similarly, using Lemma C.1 we get $\left|({\mathscr{K}}^{h}_{\phi}-{\mathscr{K}}_{\phi})(x,i;y,j)\right|\leq c^{\prime}{\mathrm{e}}^{-\xi^{\prime}(\|x-y\|_{1}+\mathrm{dist}{x,l_{R}})}$. Together with the bound (C.14) we obtain
$\left|({\mathscr{G}}_{\phi}^{(\mu)}-{\mathscr{K}}_{\phi})(x,i;y,j)\right|\leq c^{\prime\prime}{\mathrm{e}}^{-\xi^{\prime\prime}(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}.$ (C.17)
Taking the minimum of (C.16) and (C.17) yields the second bound of the Lemma. ∎
C.3 Comparison of quasi-adiabatic and spectral evolutions
Remaining in the setting of the previous subsection, we consider a closed interval $[\phi^{(i)},\phi^{(f)}]\subset{\mathbb{R}}$ of fluxes where $\mu$ is in a spectral gap $\Delta^{\prime}$ of ${\mathscr{H}}_{\phi}$. For any $\phi\in[\phi^{(i)},\phi^{(f)}]$ we let ${\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}$ be the unique solution to
${\mathscr{U}}^{\mathrm{qa}}_{(\phi^{(i)},\phi^{(i)})}=\operatorname{\mathbbm{1}},\quad{\mathrm{i}}\partial_{\phi}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}={\mathscr{K}}_{\phi}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}.$ (C.18)
Similarly, let ${\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})}$ be the unique solution to
${\mathscr{U}}^{(\mu)}_{(\phi^{(i)},\phi^{(i)})}=\operatorname{\mathbbm{1}},\quad{\mathrm{i}}\partial_{\phi}{\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})}={\mathscr{G}}^{(\mu)}_{\phi}{\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})}.$ (C.19)
Both unitary evolutions are generated by exponentially local generators, so that the following lemma applies to both:
Lemma C.4.
Let $[\phi^{(i)},\phi^{(f)}]\ni\phi\mapsto{\mathscr{A}}_{\phi}$ be a family of bounded self adjoint operators and suppose there are constants $c,\xi>0$ such that
$\left|{\mathscr{A}}_{\phi}(x,i;y,j)\right|\leq c{\mathrm{e}}^{-2\xi\|x-y\|_{1}}$ (C.20)
@@ C.3 Comparison of quasi-adiabatic and spectral evolutions
for all $\phi\in[\phi^{(i)},\phi^{(f)}]$. Let ${\mathscr{V}}_{\phi}$ be the unique solution to ${\mathscr{V}}_{\phi^{(i)}}=\operatorname{\mathbbm{1}}$ and ${\mathrm{i}}\partial_{\phi}{\mathscr{V}}_{\phi}={\mathscr{A}}_{\phi}{\mathscr{V}}_{\phi}$. Then there is a constant $C>0$ such that
$\left|{\mathscr{V}}_{\phi}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\xi\|x-y\|_{1}}.$ (C.21)
@@
Define
$a_{n}(x,y):=\sup_{\begin{subarray}{c}\vec{\phi}\in[\phi^{(i)},\phi^{(f)}]^{n}\\ i,j\in\{1,\cdots,2m\}\end{subarray}}\,\left|({\mathscr{A}}_{\phi_{1}}\cdots{\mathscr{A}}_{\phi_{n}})(x,i;y,j)\right|.$ (C.22)
for all $n\geq 1$. We show by induction that
$a_{n}(x,y)\leq c^{n}\kappa^{n-1}{\mathrm{e}}^{-\xi\|x-y\|_{1}}$ (C.23)
where
$\kappa=2m\,\sup_{x\in{\mathbb{Z}}^{2}}\,\sum_{z\in{\mathbb{Z}}^{2}}{\mathrm{e}}^{-\xi\|x-z\|_{1}}.$ (C.24)
By the assumption (C.20) the claimed bound holds true for $n=1$. Proceeding by induction,
$a_{n+1}(x,y)$ $\leq\sup_{\phi\in[\phi_{1},\phi_{2}]}\,\sup_{i\in\{1,\cdots,2m\}}\,\sum_{z,i^{\prime}}\,\left|{\mathscr{A}}_{\phi}(x,i;z,i^{\prime})\right|\times a_{n}(z,y)$ (C.25)
$\leq c^{n+1}\kappa^{n-1}\sum_{z,i^{\prime}}{\mathrm{e}}^{-2\xi\|x-z\|_{1}}\times{\mathrm{e}}^{-\xi\|z-y\|_{1}}\leq c^{n+1}\kappa^{n}{\mathrm{e}}^{-\xi\|x-y\|_{1}}$ (C.26)
as required.
The bound on the matrix elements of ${\mathscr{V}}_{\phi}$ now follows from
$\left|({\mathscr{V}}_{\phi}-\operatorname{\mathbbm{1}})(x,i;y,j)\right|\leq\sum_{n=1}^{\infty}\frac{\left|\phi\right|^{n}}{n!}a_{n}(x,y).$ (C.27)
Lemma C.5.
The difference ${\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}-{\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})}$ is trace class for all $\phi\in[\phi^{(i)},\phi^{(f)}]$.
@@
It is sufficient to show that $({\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})})^{*}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}-\operatorname{\mathbbm{1}}$ is trace class. We have
$({\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})})^{*}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}-\operatorname{\mathbbm{1}}={\mathrm{i}}\int_{\phi^{(i)}}^{\phi}{\mathrm{d}}\phi^{\prime}({\mathscr{U}}^{(\mu)}_{(\phi^{\prime},\phi^{(i)})})^{*}\,({\mathscr{G}}^{(\mu)}_{\phi^{\prime}}-{\mathscr{K}}_{\phi^{\prime}})\,{\mathscr{U}}^{\mathrm{qa}}_{(\phi^{\prime},\phi^{(i)})}.$ (C.28)
By Lemma C.4 and the exponential locality of ${\mathscr{G}}_{\phi}^{(\mu)}$ and ${\mathscr{K}}_{\phi}$, the unitaries ${\mathscr{U}}^{(\mu)}_{(\phi^{\prime},\phi^{(i)})}$ and ${\mathscr{U}}^{\mathrm{qa}}_{(\phi^{\prime},\phi^{(i)})}$ are exponentially local. Combining this with the bound on ${\mathscr{G}}^{(\mu)}_{\phi}-{\mathscr{K}}_{\phi}$ from Lemma C.3 we find that there are constants $c,\xi>0$ such that
$\left|\big{(}({\mathscr{U}}^{(\mu)}_{(\phi,\phi^{(i)})})^{*}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}-\operatorname{\mathbbm{1}}\big{)}(x,i;y,j)\right|\leq c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\|x\|_{1})}.$ (C.29)
The trace class claim now follows from Lemma 1 of [1]. ∎
Proposition C.6.
For all $\phi\in[\phi^{(i)},\phi^{(f)}]$, the difference of projections
${\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}{\mathscr{P}}_{\phi_{1}}^{(\mu)}{\mathscr{U}}^{\mathrm{qa}}_{(\phi,\phi^{(i)})}-{\mathscr{P}}^{(\mu)}_{\phi}$
is trace class, and the trace vanishes.
Proof.